# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-4/describe-the-difference-between-related-words 52 89 88 100 100
grade-4/to-have-use-the-correct-form 63 100 100 100 100
grade-4/use-subordinating-conjunctions 49 100 64 100 100
grade-4/use-the-correct-homophone 39 100 100 100 100
grade-4/is-it-a-complete-sentence-or-a-fragment 67 70 97 100 100
grade-4/choose-the-best-concluding-sentence 46 100 100 100 100
grade-4/identify-the-narrative-point-of-view 52 88 76 82 100
grade-4/formatting-and-capitalizing-titles 42 50 46 62 54
grade-4/choose-the-best-topic-sentence 20 100 100 100 100
grade-4/determine-the-meaning-of-idioms-from-context-set-1 48 100 100 100 100
grade-4/punctuating-dialogue 26 92 78 92 81
grade-4/determine-the-meanings-of-similes-and-metaphors 51 100 100 100 100
grade-4/identify-similes-and-metaphors 30 79 76 100 100
grade-4/use-the-correct-modal-verb 65 95 94 100 100
grade-4/determine-the-meaning-of-idioms-from-context-set-2 55 100 100 100 100
grade-4/identify-the-meaning-of-idioms-and-adages-set-2 54 100 100 100 100
grade-4/use-the-perfect-verb-tenses 84 100 92 100 100
grade-4/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 53 93 100 100 82
grade-4/use-the-progressive-verb-tenses 75 84 82 100 100
grade-4/what-does-the-modal-verb-show 47 73 70 100 100
grade-4/identify-the-meaning-of-idioms-and-adages-set-1 49 75 100 100 100
grade-4/use-thesaurus-entries 42 100 100 94 100
@@
grade-5/determine-the-meanings-of-words-with-greek-and-latin-roots 55 100 100 100 100
grade-5/describe-the-difference-between-related-words 59 100 88 100 100
grade-5/use-the-correct-homophone 51 100 100 100 100
grade-5/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 21 100 100 100 100
grade-5/analogies 67 72 73 100 100
grade-5/is-it-a-complete-sentence-or-a-fragment 53 76 76 100 100
grade-5/choose-reasons-to-support-an-opinion 66 64 91 92 91
grade-5/choose-the-synonym 47 100 100 100 100
grade-5/select-the-words-that-dont-belong 33 66 67 66 65
grade-5/is-it-a-complete-sentence-or-a-run-on 63 70 65 100 100
grade-5/use-academic-vocabulary-in-context 61 100 100 100 100
grade-5/use-adjectives-to-compare 63 68 100 100 84
grade-5/capitalizing-titles 57 61 44 97 81
grade-5/use-the-correct-subject-or-verb 68 76 69 71 100
grade-5/choose-between-subject-and-object-personal-pronouns 38 85 74 100 100
grade-5/choose-between-personal-and-reflexive-pronouns 53 85 92 100 100
grade-5/choose-the-best-concluding-sentence 39 100 100 100 100
grade-5/create-compound-sentences 37 46 22 72 86
grade-5/is-it-a-complete-sentence-a-fragment-or-a-run-on 59 87 69 100 100
grade-5/is-the-word-an-adjective-or-adverb 69 100 92 100 100
grade-5/identify-the-narrative-point-of-view 60 78 73 75 94
grade-5/use-actions-and-dialogue-to-understand-characters 46 100 100 100 100
grade-5/use-adverbs-to-compare 84 64 63 97 77
grade-5/is-the-sentence-in-the-past-present-or-future-tense 44 87 60 100 100
grade-5/formatting-and-capitalizing-titles 50 46 46 46 37
grade-5/which-definition-matches-the-sentence 57 100 90 100 100
grade-5/choose-the-antonym 71 100 100 100 100
grade-5/choose-the-best-topic-sentence 48 100 100 100 100
grade-5/read-poetry 54 100 90 100 100
grade-5/show-character-emotions-and-traits 54 70 91 91 92
grade-5/identify-text-structures 71 90 100 100 100
grade-5/determine-the-meaning-of-idioms-from-context-set-1 46 100 100 100 100
grade-5/punctuating-dialogue 42 87 81 84 84
grade-5/determine-the-meanings-of-similes-and-metaphors 61 100 100 100 100
grade-5/use-key-details-to-determine-the-main-idea 27 100 100 100 100
grade-5/choose-between-adjectives-and-adverbs 49 100 80 100 83
grade-5/use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots 42 100 100 100 100
grade-5/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 28 100 100 100 100
grade-5/use-coordinating-conjunctions 48 100 47 100 100
grade-5/identify-similes-and-metaphors 73 71 72 100 100
grade-5/interpret-the-meaning-of-an-allusion-from-its-source 68 100 100 100 100
grade-5/positive-and-negative-connotation 75 100 67 100 100
grade-5/distinguish-facts-from-opinions 72 100 100 100 100
grade-5/choose-the-best-transition 38 100 86 100 100
grade-5/draw-inferences-from-a-text 41 100 79 83 94
grade-5/is-the-sentence-simple-or-compound 64 77 67 100 100
grade-5/identify-elements-of-poetry 42 57 45 48 88
grade-5/identify-plurals-singular-possessives-and-plural-possessives 60 100 56 100 100
@@
grade-5/determine-the-meanings-of-greek-and-latin-roots 41 100 70 100 100
grade-5/analyze-the-effects-of-figures-of-speech-on-meaning-and-tone 56 100 100 100 100
grade-5/words-with-sub 56 63 83 100 100
grade-5/words-with-pre 40 100 100 100 100
grade-5/use-the-correct-modal-verb 54 100 97 100 100
grade-5/use-the-perfect-verb-tenses 35 100 77 100 100
grade-5/words-with-less 67 100 100 100 100
grade-5/words-with-un-dis-in-im-and-non 50 100 100 100 100
grade-5/determine-the-meaning-of-idioms-from-context-set-2 53 100 100 100 100
grade-5/choose-between-the-past-tense-and-past-participle 72 73 100 100 100
grade-5/words-with-re 64 100 100 100 100
grade-5/use-the-correct-pair-of-correlative-conjunctions 64 64 41 76 84
grade-5/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 37 100 100 100 92
grade-5/what-does-the-modal-verb-show 57 98 76 100 100
grade-5/use-the-progressive-verb-tenses 73 80 92 100 100
grade-5/words-with-ful 77 100 100 100 100
grade-5/compound-subjects-and-objects-with-personal-pronouns 84 72 55 94 100
grade-5/words-with-able-and-ible 54 100 100 100 100
grade-5/determine-the-themes-of-short-stories 57 84 69 73 100
grade-5/use-the-correct-frequently-confused-word 71 56 73 75 75
grade-5/words-with-mis 39 100 77 100 100
grade-5/which-sentence-is-more-formal 75 100 100 100 100
grade-5/is-the-sentence-simple-compound-or-complex 60 40 55 95 98
grade-6/determine-the-meanings-of-words-with-greek-and-latin-roots 40 100 100 100 100
grade-6/describe-the-difference-between-related-words 52 100 100 100 100
grade-6/use-the-correct-homophone 53 97 100 100 100
grade-6/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 49 100 100 100 100
grade-6/analogies 43 83 64 100 100
grade-6/is-it-a-complete-sentence-or-a-fragment 29 77 53 100 100
grade-6/choose-between-subject-and-object-pronouns 43 94 65 100 100
grade-6/analogies-challenge 19 80 75 87 100
grade-6/is-it-a-complete-sentence-or-a-run-on 59 70 64 100 100
grade-6/transitions-with-conjunctive-adverbs 69 71 73 87 68
grade-6/determine-the-authors-point-of-view 66 66 100 100 100
grade-6/compare-two-texts-with-different-genres 37 100 100 100 100
grade-6/choose-evidence-to-support-a-claim 60 45 86 78 86
grade-6/use-personification 53 80 51 51 46
grade-6/is-the-word-an-adjective-or-adverb 59 76 100 100 100
grade-6/is-it-a-complete-sentence-a-fragment-or-a-run-on 52 82 72 100 100
grade-6/which-definition-matches-the-sentence 36 100 19 100 100
grade-6/identify-counterclaims 53 80 100 100 100
grade-6/identify-text-structures 74 72 100 100 100
grade-6/use-the-correct-pair-of-correlative-conjunctions 60 44 50 56 77
@@
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-6/is-the-pronoun-reflexive-or-intensive 47 56 46 63 74
grade-6/interpret-figures-of-speech 61 100 100 100 100
grade-6/classify-figures-of-speech 76 100 100 81 80
grade-6/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 31 73 79 100 100
grade-6/vocabulary-review-trace-an-argument 40 100 100 100 100
grade-6/identify-the-authors-purpose 39 100 100 100 100
grade-6/is-it-a-direct-object-or-an-indirect-object 44 49 71 70 100
grade-6/compound-subjects-and-objects-with-pronouns 71 74 48 100 100
grade-6/identify-vague-pronoun-references 82 66 81 100 76
grade-6/choose-between-adjectives-and-adverbs 34 100 100 100 100
grade-6/what-does-the-punctuation-suggest 46 49 60 79 72
grade-6/suggest-appropriate-revisions 55 67 83 91 73
grade-6/classify-logical-fallacies 58 69 70 100 100
grade-6/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 40 100 100 100 100
grade-6/use-the-correct-frequently-confused-word 42 81 79 69 78
grade-6/distinguish-facts-from-opinions 43 100 100 100 100
grade-6/use-the-correct-verb-with-compound-subjects 64 74 72 91 100
grade-6/determine-the-main-idea-of-a-passage 46 100 90 86 100
grade-6/analyze-short-stories 35 64 80 63 71
grade-6/identify-plurals-singular-possessives-and-plural-possessives 43 75 69 100 100
grade-6/which-sentence-is-more-formal 99 100 100 100 100
grade-6/determine-the-meanings-of-greek-and-latin-roots 36 100 96 100 100
@@
grade-6/analyze-passages-from-roll-of-thunder-hear-my-cry-part-1 67 100 94 100 100
grade-6/use-the-correct-modal-verb 64 100 100 100 100
grade-6/use-the-perfect-verb-tenses 63 92 86 82 100
grade-6/determine-the-meaning-of-idioms-from-context-set-2 72 100 100 100 100
grade-6/choose-between-the-past-tense-and-past-participle 64 100 93 100 100
grade-6/use-the-correct-subject-or-verb 68 66 59 73 79
grade-6/evaluate-newspaper-headlines-for-bias 36 71 80 100 84
grade-6/determine-the-meaning-of-idioms-from-context-set-1 44 100 100 100 100
grade-6/use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots 16 100 97 100 100
grade-6/what-does-the-modal-verb-show 65 100 92 100 100
grade-6/use-the-progressive-verb-tenses 71 100 85 100 100
grade-6/use-coordinating-conjunctions 73 84 58 100 100
grade-6/is-the-sentence-simple-compound-or-complex 39 43 51 75 100
@@
grade-7/determine-the-meanings-of-words-with-greek-and-latin-roots 43 100 19 100 100
grade-7/describe-the-difference-between-related-words 60 100 100 100 100
grade-7/use-the-correct-homophone 44 100 100 100 100
grade-7/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 27 100 82 100 100
grade-7/analogies 51 100 71 100 100
grade-7/is-it-a-complete-sentence-or-a-fragment 67 74 67 100 100
grade-7/is-it-a-phrase-or-a-clause 78 59 42 98 74
grade-7/choose-between-subject-and-object-pronouns 69 100 71 100 100
grade-7/is-the-sentence-simple-compound-complex-or-compound-complex 28 19 25 34 70
grade-7/analogies-challenge 48 73 81 99 100
grade-7/is-it-a-complete-sentence-or-a-run-on 64 84 60 100 100
grade-7/are-the-modifiers-used-correctly 80 74 80 64 68
grade-7/recall-the-source-of-an-allusion 57 100 93 100 100
grade-7/transitions-with-conjunctive-adverbs 54 83 51 70 82
grade-7/determine-the-authors-point-of-view 65 54 100 100 100
grade-7/choose-the-synonym 54 100 100 100 100
grade-7/choose-evidence-to-support-a-claim 45 46 86 55 86
grade-7/use-personification 45 43 46 78 71
grade-7/identify-plagiarism 36 44 37 41 23
grade-7/is-the-word-an-adjective-or-adverb 59 84 89 100 100
grade-7/is-it-a-complete-sentence-a-fragment-or-a-run-on 59 69 64 100 100
grade-7/which-definition-matches-the-sentence 58 100 46 100 100
grade-7/choose-the-antonym 60 100 92 100 100
grade-7/identify-text-structures 58 69 100 100 100
grade-7/use-the-correct-pair-of-correlative-conjunctions 64 64 55 79 82
grade-7/is-the-pronoun-reflexive-or-intensive 33 48 70 73 68
grade-7/interpret-figures-of-speech 46 100 100 100 100
grade-7/classify-figures-of-speech 64 100 100 88 69
grade-7/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 32 83 76 100 100
grade-7/vocabulary-review-trace-an-argument 25 82 100 100 100
grade-7/identify-the-authors-purpose 50 100 100 100 100
grade-7/decide-whether-ellipses-are-used-appropriately 63 47 45 76 68
grade-7/compound-subjects-and-objects-with-pronouns 47 78 55 100 100
grade-7/identify-vague-pronoun-references 45 67 83 83 72
grade-7/choose-between-adjectives-and-adverbs 65 92 92 96 100
grade-7/what-does-the-punctuation-suggest 46 59 68 77 72
grade-7/suggest-appropriate-revisions 61 82 59 77 82
grade-7/determine-the-themes-of-short-stories 43 61 80 63 70
grade-7/classify-logical-fallacies 55 72 82 100 93
grade-7/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 57 100 100 100 100
grade-7/interpret-the-meaning-of-an-allusion-from-its-source 50 100 100 100 100
grade-7/use-the-correct-frequently-confused-word 54 76 70 79 75
grade-7/distinguish-facts-from-opinions 40 100 91 100 100
grade-7/use-the-correct-verb-with-compound-subjects 64 56 62 83 100
grade-7/determine-the-main-idea-of-a-passage 46 77 100 100 100
grade-7/identify-plurals-singular-possessives-and-plural-possessives 45 82 54 100 100
grade-7/which-sentence-is-more-formal 57 100 100 100 100
grade-7/determine-the-meanings-of-greek-and-latin-roots 23 100 74 100 100
grade-7/use-parallel-structure 43 82 57 74 95
@@
grade-8/determine-the-meanings-of-words-with-greek-and-latin-roots 25 100 100 100 100
grade-8/describe-the-difference-between-related-words 69 100 82 100 100
grade-8/use-the-correct-homophone 57 82 99 100 100
grade-8/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 24 100 67 100 100
grade-8/analogies 62 87 64 100 95
grade-8/is-it-a-complete-sentence-or-a-fragment 64 100 68 100 100
grade-8/choose-between-subject-and-object-pronouns 25 77 56 100 100
grade-8/is-the-sentence-simple-compound-complex-or-compound-complex 16 25 25 52 63
grade-8/analogies-challenge 32 94 57 82 100
grade-8/is-it-a-complete-sentence-or-a-run-on 59 65 71 100 100
grade-8/are-the-modifiers-used-correctly 58 59 72 85 75
grade-8/recall-the-source-of-an-allusion 53 100 100 100 100
grade-8/transitions-with-conjunctive-adverbs 79 82 68 84 79
grade-8/determine-the-authors-point-of-view 64 43 100 100 100
grade-8/choose-the-synonym 46 100 100 100 100
grade-8/compare-two-texts-with-different-genres 41 86 100 100 100
grade-8/choose-evidence-to-support-a-claim 26 46 86 64 86
grade-8/choose-between-personal-and-reflexive-pronouns 37 79 72 100 100
grade-8/use-personification 54 90 43 100 83
grade-8/identify-plagiarism 65 49 42 45 48
grade-8/identify-the-narrative-point-of-view 56 60 63 85 100
grade-8/is-the-word-an-adjective-or-adverb 50 71 100 100 100
grade-8/identify-counterclaims 49 80 100 100 100
grade-8/is-it-a-complete-sentence-a-fragment-or-a-run-on 37 67 73 100 100
grade-8/which-definition-matches-the-sentence 46 100 9 100 100
grade-8/choose-the-antonym 63 100 83 100 100
grade-8/identify-text-structures 45 63 100 100 100
grade-8/use-the-correct-pair-of-correlative-conjunctions 54 38 64 77 74
grade-8/identify-active-and-passive-voice 64 70 47 87 100

### v2

@@ Appendix D Summary of Skills
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-4/describe-the-difference-between-related-words
52
89
88
100
100
grade-4/to-have-use-the-correct-form
63
100
100
100
100
grade-4/use-subordinating-conjunctions
49
100
64
100
100
grade-4/use-the-correct-homophone
39
100
100
100
100
grade-4/is-it-a-complete-sentence-or-a-fragment
67
70
97
100
100
grade-4/choose-the-best-concluding-sentence
46
100
100
100
100
grade-4/identify-the-narrative-point-of-view
52
88
76
82
100
grade-4/formatting-and-capitalizing-titles
42
50
46
62
54
grade-4/choose-the-best-topic-sentence
20
100
100
100
100
grade-4/determine-the-meaning-of-idioms-from-context-set-1
48
100
100
100
100
grade-4/punctuating-dialogue
26
92
78
92
81
grade-4/determine-the-meanings-of-similes-and-metaphors
51
100
100
100
100
grade-4/identify-similes-and-metaphors
30
79
76
100
100
grade-4/use-the-correct-modal-verb
65
95
94
100
100
grade-4/determine-the-meaning-of-idioms-from-context-set-2
55
100
100
100
100
grade-4/identify-the-meaning-of-idioms-and-adages-set-2
54
100
100
100
100
grade-4/use-the-perfect-verb-tenses
84
100
92
100
100
grade-4/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
53
93
100
100
82
grade-4/use-the-progressive-verb-tenses
75
84
82
100
100
grade-4/what-does-the-modal-verb-show
47
73
70
100
100
grade-4/identify-the-meaning-of-idioms-and-adages-set-1
49
75
100
100
100
grade-4/use-thesaurus-entries
42
100
100
94
100
@@
grade-5/determine-the-meanings-of-words-with-greek-and-latin-roots
55
100
100
100
100
grade-5/describe-the-difference-between-related-words
59
100
88
100
100
grade-5/use-the-correct-homophone
51
100
100
100
100
grade-5/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
21
100
100
100
100
grade-5/analogies
67
72
73
100
100
grade-5/is-it-a-complete-sentence-or-a-fragment
53
76
76
100
100
grade-5/choose-reasons-to-support-an-opinion
66
64
91
92
91
grade-5/choose-the-synonym
47
100
100
100
100
grade-5/select-the-words-that-dont-belong
33
66
67
66
65
grade-5/is-it-a-complete-sentence-or-a-run-on
63
70
65
100
100
grade-5/use-academic-vocabulary-in-context
61
100
100
100
100
grade-5/use-adjectives-to-compare
63
68
100
100
84
grade-5/capitalizing-titles
57
61
44
97
81
grade-5/use-the-correct-subject-or-verb
68
76
69
71
100
grade-5/choose-between-subject-and-object-personal-pronouns
38
85
74
100
100
grade-5/choose-between-personal-and-reflexive-pronouns
53
85
92
100
100
grade-5/choose-the-best-concluding-sentence
39
100
100
100
100
grade-5/create-compound-sentences
37
46
22
72
86
grade-5/is-it-a-complete-sentence-a-fragment-or-a-run-on
59
87
69
100
100
grade-5/is-the-word-an-adjective-or-adverb
69
100
92
100
100
grade-5/identify-the-narrative-point-of-view
60
78
73
75
94
grade-5/use-actions-and-dialogue-to-understand-characters
46
100
100
100
100
grade-5/use-adverbs-to-compare
84
64
63
97
77
grade-5/is-the-sentence-in-the-past-present-or-future-tense
44
87
60
100
100
grade-5/formatting-and-capitalizing-titles
50
46
46
46
37
grade-5/which-definition-matches-the-sentence
57
100
90
100
100
grade-5/choose-the-antonym
71
100
100
100
100
grade-5/choose-the-best-topic-sentence
48
100
100
100
100
grade-5/read-poetry
54
100
90
100
100
grade-5/show-character-emotions-and-traits
54
70
91
91
92
grade-5/identify-text-structures
71
90
100
100
100
grade-5/determine-the-meaning-of-idioms-from-context-set-1
46
100
100
100
100
grade-5/punctuating-dialogue
42
87
81
84
84
grade-5/determine-the-meanings-of-similes-and-metaphors
61
100
100
100
100
grade-5/use-key-details-to-determine-the-main-idea
27
100
100
100
100
grade-5/choose-between-adjectives-and-adverbs
49
100
80
100
83
grade-5/use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots
42
100
100
100
100
grade-5/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
28
100
100
100
100
grade-5/use-coordinating-conjunctions
48
100
47
100
100
grade-5/identify-similes-and-metaphors
73
71
72
100
100
grade-5/interpret-the-meaning-of-an-allusion-from-its-source
68
100
100
100
100
grade-5/positive-and-negative-connotation
75
100
67
100
100
grade-5/distinguish-facts-from-opinions
72
100
100
100
100
grade-5/choose-the-best-transition
38
100
86
100
100
grade-5/draw-inferences-from-a-text
41
100
79
83
94
grade-5/is-the-sentence-simple-or-compound
64
77
67
100
100
grade-5/identify-elements-of-poetry
42
57
45
48
88
grade-5/identify-plurals-singular-possessives-and-plural-possessives
60
100
56
100
100
grade-5/determine-the-meanings-of-greek-and-latin-roots
41
100
70
100
100
grade-5/analyze-the-effects-of-figures-of-speech-on-meaning-and-tone
56
100
100
100
100
grade-5/words-with-sub
56
63
83
100
100
grade-5/words-with-pre
40
100
100
100
100
grade-5/use-the-correct-modal-verb
54
100
97
100
100
grade-5/use-the-perfect-verb-tenses
35
100
77
100
100
grade-5/words-with-less
67
100
100
100
100
grade-5/words-with-un-dis-in-im-and-non
50
100
100
100
100
grade-5/determine-the-meaning-of-idioms-from-context-set-2
53
100
100
100
100
grade-5/choose-between-the-past-tense-and-past-participle
72
73
100
100
100
grade-5/words-with-re
64
100
100
100
100
grade-5/use-the-correct-pair-of-correlative-conjunctions
64
64
41
76
84
grade-5/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
37
100
100
100
92
grade-5/what-does-the-modal-verb-show
57
98
76
100
100
grade-5/use-the-progressive-verb-tenses
73
80
92
100
100
grade-5/words-with-ful
77
100
100
100
100
grade-5/compound-subjects-and-objects-with-personal-pronouns
84
72
55
94
100
grade-5/words-with-able-and-ible
54
100
100
100
100
grade-5/determine-the-themes-of-short-stories
57
84
69
73
100
grade-5/use-the-correct-frequently-confused-word
71
56
73
75
75
grade-5/words-with-mis
39
100
77
100
100
grade-5/which-sentence-is-more-formal
75
100
100
100
100
grade-5/is-the-sentence-simple-compound-or-complex
60
40
55
95
98
@@
grade-6/determine-the-meanings-of-words-with-greek-and-latin-roots
40
100
100
100
100
grade-6/describe-the-difference-between-related-words
52
100
100
100
100
grade-6/use-the-correct-homophone
53
97
100
100
100
grade-6/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
49
100
100
100
100
grade-6/analogies
43
83
64
100
100
grade-6/is-it-a-complete-sentence-or-a-fragment
29
77
53
100
100
grade-6/choose-between-subject-and-object-pronouns
43
94
65
100
100
grade-6/analogies-challenge
19
80
75
87
100
grade-6/is-it-a-complete-sentence-or-a-run-on
59
70
64
100
100
grade-6/transitions-with-conjunctive-adverbs
69
71
73
87
68
grade-6/determine-the-authors-point-of-view
66
66
100
100
100
grade-6/compare-two-texts-with-different-genres
37
100
100
100
100
grade-6/choose-evidence-to-support-a-claim
60
45
86
78
86
grade-6/use-personification
53
80
51
51
46
grade-6/is-the-word-an-adjective-or-adverb
59
76
100
100
100
grade-6/is-it-a-complete-sentence-a-fragment-or-a-run-on
52
82
72
100
100
grade-6/which-definition-matches-the-sentence
36
100
19
100
100
grade-6/identify-counterclaims
53
80
100
100
100
grade-6/identify-text-structures
74
72
100
100
100
grade-6/use-the-correct-pair-of-correlative-conjunctions
60
44
50
56
77
@@
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-6/is-the-pronoun-reflexive-or-intensive
47
56
46
63
74
grade-6/interpret-figures-of-speech
61
100
100
100
100
grade-6/classify-figures-of-speech
76
100
100
81
80
grade-6/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
31
73
79
100
100
grade-6/vocabulary-review-trace-an-argument
40
100
100
100
100
grade-6/identify-the-authors-purpose
39
100
100
100
100
grade-6/is-it-a-direct-object-or-an-indirect-object
44
49
71
70
100
grade-6/compound-subjects-and-objects-with-pronouns
71
74
48
100
100
grade-6/identify-vague-pronoun-references
82
66
81
100
76
grade-6/choose-between-adjectives-and-adverbs
34
100
100
100
100
grade-6/what-does-the-punctuation-suggest
46
49
60
79
72
grade-6/suggest-appropriate-revisions
55
67
83
91
73
grade-6/classify-logical-fallacies
58
69
70
100
100
grade-6/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
40
100
100
100
100
grade-6/use-the-correct-frequently-confused-word
42
81
79
69
78
grade-6/distinguish-facts-from-opinions
43
100
100
100
100
grade-6/use-the-correct-verb-with-compound-subjects
64
74
72
91
100
grade-6/determine-the-main-idea-of-a-passage
46
100
90
86
100
grade-6/analyze-short-stories
35
64
80
63
71
grade-6/identify-plurals-singular-possessives-and-plural-possessives
43
75
69
100
100
grade-6/which-sentence-is-more-formal
99
100
100
100
100
grade-6/determine-the-meanings-of-greek-and-latin-roots
36
100
96
100
100
@@
grade-6/analyze-passages-from-roll-of-thunder-hear-my-cry-part-1
67
100
94
100
100
grade-6/use-the-correct-modal-verb
64
100
100
100
100
grade-6/use-the-perfect-verb-tenses
63
92
86
82
100
grade-6/determine-the-meaning-of-idioms-from-context-set-2
72
100
100
100
100
grade-6/choose-between-the-past-tense-and-past-participle
64
100
93
100
100
grade-6/use-the-correct-subject-or-verb
68
66
59
73
79
grade-6/evaluate-newspaper-headlines-for-bias
36
71
80
100
84
grade-6/determine-the-meaning-of-idioms-from-context-set-1
44
100
100
100
100
grade-6/use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots
16
100
97
100
100
grade-6/what-does-the-modal-verb-show
65
100
92
100
100
grade-6/use-the-progressive-verb-tenses
71
100
85
100
100
grade-6/use-coordinating-conjunctions
73
84
58
100
100
grade-6/is-the-sentence-simple-compound-or-complex
39
43
51
75
100
@@
grade-7/determine-the-meanings-of-words-with-greek-and-latin-roots
43
100
19
100
100
grade-7/describe-the-difference-between-related-words
60
100
100
100
100
grade-7/use-the-correct-homophone
44
100
100
100
100
grade-7/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
27
100
82
100
100
grade-7/analogies
51
100
71
100
100
grade-7/is-it-a-complete-sentence-or-a-fragment
67
74
67
100
100
grade-7/is-it-a-phrase-or-a-clause
78
59
42
98
74
grade-7/choose-between-subject-and-object-pronouns
69
100
71
100
100
grade-7/is-the-sentence-simple-compound-complex-or-compound-complex
28
19
25
34
70
grade-7/analogies-challenge
48
73
81
99
100
grade-7/is-it-a-complete-sentence-or-a-run-on
64
84
60
100
100
grade-7/are-the-modifiers-used-correctly
80
74
80
64
68
grade-7/recall-the-source-of-an-allusion
57
100
93
100
100
grade-7/transitions-with-conjunctive-adverbs
54
83
51
70
82
grade-7/determine-the-authors-point-of-view
65
54
100
100
100
grade-7/choose-the-synonym
54
100
100
100
100
grade-7/choose-evidence-to-support-a-claim
45
46
86
55
86
grade-7/use-personification
45
43
46
78
71
grade-7/identify-plagiarism
36
44
37
41
23
grade-7/is-the-word-an-adjective-or-adverb
59
84
89
100
100
grade-7/is-it-a-complete-sentence-a-fragment-or-a-run-on
59
69
64
100
100
grade-7/which-definition-matches-the-sentence
58
100
46
100
100
grade-7/choose-the-antonym
60
100
92
100
100
grade-7/identify-text-structures
58
69
100
100
100
grade-7/use-the-correct-pair-of-correlative-conjunctions
64
64
55
79
82
grade-7/is-the-pronoun-reflexive-or-intensive
33
48
70
73
68
grade-7/interpret-figures-of-speech
46
100
100
100
100
grade-7/classify-figures-of-speech
64
100
100
88
69
grade-7/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
32
83
76
100
100
grade-7/vocabulary-review-trace-an-argument
25
82
100
100
100
grade-7/identify-the-authors-purpose
50
100
100
100
100
grade-7/decide-whether-ellipses-are-used-appropriately
63
47
45
76
68
grade-7/compound-subjects-and-objects-with-pronouns
47
78
55
100
100
grade-7/identify-vague-pronoun-references
45
67
83
83
72
grade-7/choose-between-adjectives-and-adverbs
65
92
92
96
100
grade-7/what-does-the-punctuation-suggest
46
59
68
77
72
grade-7/suggest-appropriate-revisions
61
82
59
77
82
grade-7/determine-the-themes-of-short-stories
43
61
80
63
70
grade-7/classify-logical-fallacies
55
72
82
100
93
grade-7/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
57
100
100
100
100
grade-7/interpret-the-meaning-of-an-allusion-from-its-source
50
100
100
100
100
grade-7/use-the-correct-frequently-confused-word
54
76
70
79
75
grade-7/distinguish-facts-from-opinions
40
100
91
100
100
grade-7/use-the-correct-verb-with-compound-subjects
64
56
62
83
100
grade-7/determine-the-main-idea-of-a-passage
46
77
100
100
100
grade-7/identify-plurals-singular-possessives-and-plural-possessives
45
82
54
100
100
grade-7/which-sentence-is-more-formal
57
100
100
100
100
grade-7/determine-the-meanings-of-greek-and-latin-roots
23
100
74
100
100
grade-7/use-parallel-structure
43
82
57
74
95
@@
grade-8/determine-the-meanings-of-words-with-greek-and-latin-roots
25
100
100
100
100
grade-8/describe-the-difference-between-related-words
69
100
82
100
100
grade-8/use-the-correct-homophone
57
82
99
100
100
grade-8/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
24
100
67
100
100
grade-8/analogies
62
87
64
100
95
grade-8/is-it-a-complete-sentence-or-a-fragment
64
100
68
100
100
grade-8/choose-between-subject-and-object-pronouns
25
77
56
100
100
grade-8/is-the-sentence-simple-compound-complex-or-compound-complex
16
25
25
52
63
grade-8/analogies-challenge
32
94
57
82
100
grade-8/is-it-a-complete-sentence-or-a-run-on
59
65
71
100
100
grade-8/are-the-modifiers-used-correctly
58
59
72
85
75
grade-8/recall-the-source-of-an-allusion
53
100
100
100
100
grade-8/transitions-with-conjunctive-adverbs
79
82
68
84
79
grade-8/determine-the-authors-point-of-view
64
43
100
100
100
grade-8/choose-the-synonym
46
100
100
100
100
grade-8/compare-two-texts-with-different-genres
41
86
100
100
100
grade-8/choose-evidence-to-support-a-claim
26
46
86
64
86
grade-8/choose-between-personal-and-reflexive-pronouns
37
79
72
100
100
grade-8/use-personification
54
90
43
100
83
grade-8/identify-plagiarism
65
49
42
45
48
grade-8/identify-the-narrative-point-of-view
56
60
63
85
100
grade-8/is-the-word-an-adjective-or-adverb
50
71
100
100
100
grade-8/identify-counterclaims
49
80
100
100
100
grade-8/is-it-a-complete-sentence-a-fragment-or-a-run-on
37
67
73
100
100
grade-8/which-definition-matches-the-sentence
46
100
9
100
100
grade-8/choose-the-antonym
63
100
83
100
100
grade-8/identify-text-structures
45
63
100
100
100
grade-8/use-the-correct-pair-of-correlative-conjunctions
54
38
64
77
74
grade-8/identify-active-and-passive-voice
64
70
47
87
100